OpenStudy (anonymous):

Please Help ! Radium -266 has a half-life of 1600 years. How long would it take for 1.0g of radium -266 to decay into 0.25g? I don't understand how to work this out?

OpenStudy (anonymous):

Whats half life? any idea?

OpenStudy (anonymous):

first half life(after 1600 years) = 1/2 = .5 g left after second half life (16oo+1600) means after 3200 years = .5/2 = .25 g left So after 3200 years .25 g left .

OpenStudy (anonymous):

thats right..after every half life ,,,the species becomes half of what it was

OpenStudy (anonymous):

thanks @sarah786

OpenStudy (anonymous):

yw :)